OpenStudy (anonymous):

Show that if G is an open set and F is a closed set,then G-F is an open set and F-G is a closed set.

OpenStudy (anonymous):

who are u pls?

OpenStudy (anonymous):

haha.me lah, ain. triple y, help me please.:)

OpenStudy (anonymous):

i still thinking.. sorry..

OpenStudy (anonymous):

ok. never mine.

OpenStudy (anonymous):

ye ke 3y still thinking??

OpenStudy (anonymous):

wiah, nak answer please. hukhuk. :)

OpenStudy (anonymous):

wiah 2 sapa??

OpenStudy (anonymous):

wiah tu, alawiah ke?

OpenStudy (anonymous):

ya. alawiyah la tu. hehe.

OpenStudy (anonymous):

pandainy saya

OpenStudy (anonymous):

admin la sangat. :p

OpenStudy (anonymous):

got clue la.. the intersection of an open set is open, and the intersection of a close set is close.. that what i noe.. hehe

OpenStudy (anonymous):

my frenz said use this answer The simplest way to understand this problem is to understand exactly what G\F means in the context of unions, intersections, and complements. Quite simply, A/B = A int B' (the intersection of A and not-B) Now: if F is closed, then F' is open by the definition of a closed set. Equivalently, G being open means that G' is closed. Thus, G\F = G int F' F\G = F int G' This is a countable intersection of closed sets, which must always be closed (another theorem).

OpenStudy (anonymous):

what does A/B and G/F means?

OpenStudy (anonymous):

A/B means the intersection of A and \[B^{c}\]

OpenStudy (anonymous):

same goes to G/F

OpenStudy (anonymous):

A/B or A\B??

OpenStudy (anonymous):

both also same?if write in simbol how? is it A - B?

OpenStudy (anonymous):

ya la..

OpenStudy (anonymous):

owh.ok.thanx.do u know how to solve this question? Let q(n) = 1/2 if n is prime, and -1if otherwise, and let {an\} be defined recursively by \[a _{1} = 1, a _{n+1} = a _{n} + 2^{-n} q(n), n \ge 1\] Show that {\[a _{n}\]} is a Cauchy sequence.